I NEED HELP BADLY PLEASE. THANK YOU

The following DivisionAlgorithm computes q and r for m > 0 and n > 0.  

{Input: integers m > 0 and n > 0.}

{Output: integers q and r where m.q + r = n and 0 < r < m.}

Begin

q := 0

r := n while r > m do  

q := q + 1  

r := r – m return q and r End


Revise this algorithm to work for m > 0 and n in Z. Test your revised routine using the following inputs (a) m = 5, n = 23 (b) m= 5, n = -23
Answer #1

Solution: Modified algorithm:

Begin

q:=0

r:=n

While (|r|>m ) do

if (n<0)

q=q-1

r=r+m

else

q=q+1

r=r-m

return q & r

a)For   m=5 , n=23;

     q=0 ; r=23    as 23>5 and n>0

     q=q+1=0+1=1   and r=r-m=23-5=18 (r>m):

     q=2 and r=13   (r>m)

   q=3 and r=8    (r>m)

     q=4 and r=3      (r<m)

     23=5*4+3;

b)For   m=5 , n=-23;

     q=0 ; r=-23    as |-23|= 23>5 and n<0

     q=q-1=0-1=-1   and r=r+m=-23+5=-18 (|r|>m):

   q=-2 and |r|=-13   (|r|>m)

   q=-3 and |r|=-8    (|r|>m)

   q=-4 and |r|=-3      (|r|<m)

    q=-5 and |r|=2

    q=-5 and r=2

-23=5*-5+2;